# Higher-Rank Irreducible Cartesian Tensors for Equivariant Message Passing

### v1

@@ 4.2 Equivariant message-passing based on irreducible Cartesian tensors
The following section introduces the basic operations for constructing equivariant MPNNs based on irreducible Cartesian tensors. Using their irreducible tensor products, we also demonstrate how to build equivariant two- and many-body features, crucial for modeling many-body interactions in molecular and materials systems. Particularly, we focus on MLIPs based on equivariant MPNNs and extend the state-of-the-art MACE architecture [33] to the Cartesian basis. Following the MACE architecture, we use only tensors and tensor products of even parity. We split vectors $\mathbf{r}_{uv}=\mathbf{r}_{u}-\mathbf{r}_{v}\in\mathbb{R}^{3}$ from atom $u$ to atom $v$, schematically shown in Fig. 1, into their radial and angular components (unit vectors), i.e., $r_{uv}=\lVert\mathbf{r}_{uv}\rVert_{2}\in\mathbb{R}$ and $\hat{\mathbf{r}}_{uv}=\mathbf{r}_{uv}/r_{uv}\in\mathbb{R}^{3}$, respectively. In the $t$-th message-passing layer, edges $\{u,v\}$ are embedded using a fully connected neural network $R_{kl_{1}l_{2}l_{3}}^{(t)}:\mathbb{R}\rightarrow\mathbb{R}$ with $k$ output feature channels. The radial function $R_{kl_{1}l_{2}l_{3}}^{(t)}$ takes as an input radial distances $r_{uv}$, which are embedded through Bessel functions and multiplied by a smooth polynomial function [70]. Finally, we use irreducible Cartesian tensors $\mathbf{T}_{l}(\hat{\mathbf{r}})$, similar to spherical tensors $\mathbf{Y}^{l}(\hat{\mathbf{r}})$, to embed unit vectors into the tensor space of maximal rank $l_{\mathrm{max}}$.
@@
Equivariant many-body features. The importance of many-body terms arises from the fact that the interaction between atoms changes when additional atoms are present; see Appendix A. Furthermore, many-body terms are often required to ensure the transferability of interatomic potentials [77]. Here, we construct $(\nu+1)$-body equivariant features from $\big{(}\mathbf{A}_{ukl_{\xi}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{\xi}}}$ obtained using Eqs. (4) or (5). The $\nu$-fold Cartesian tensor product, which yields $(\nu+1)$-body features represented by an irreducible Cartesian tensor of rank $L$, reads
@@
Runtime considerations. When choosing an architecture to implement MLIPs, the runtime per force and energy evaluation is crucial. The asymptotic time complexity as a function of rank $L$ is $\mathcal{O}\left(9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$ for irreducible Cartesian tensors and $\mathcal{O}\left(L^{6}\right)$ for spherical ones; see Appendix B.3 for more details. Thus, spherical tensors are more computationally efficient for $L\rightarrow\infty$. However, state-of-the-art models and physical properties typically require $L\leq 4$ [33, 78], making sub-leading terms and implementation-dependent amplitudes crucial. Based on our analysis, for $L\leq 4$, we can expect models based on irreducible Cartesian tensors to be more computationally and memory efficient than their spherical counterparts. Nonetheless, it is essential to consider the inference times for a fair comparison, which we provide in Section 5.
Equivariance of message-passing layers. We conclude this section by giving a theoretical result that ensures the equivariance of message-passing layers based on irreducible Cartesian tensors and their irreducible tensor products to actions of the orthogonal group. The proof is provided in Appendix C.
@@ 5 Experiments and results
This section presents the results for the three benchmark data sets: rMD17, 3BPA, and acetylacetone. We describe data sets and training details in Appendix D.1 and Appendix D.2, respectively.
Table 1: Energy (E) and force (F) mean absolute errors (MAEs) for the rMD17 data set. E- and F-MAE are provided in meV and meV/Å, respectively. Results are presented for models trained using $N_{\text{train}}=950$ and $N_{\text{train}}=50$ configurations randomly drawn from the data set. Further 50 configurations are used for early stopping. All values are obtained by averaging over five independent runs, with the standard deviation provided if available. Best performances, considering the standard deviation, are highlighted in bold.
@@
Table 2: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, while the remaining 50 are used for early stopping. All values for ICTP models are obtained by averaging over five independent runs. For MACE and NequIP, the results are reported for three independent runs. The standard deviation is provided if it is available. Best performances, considering the standard deviation, are highlighted in bold. Inference time and memory consumption are measured for a batch size of 100. Inference time is reported per structure in ms, while memory consumption is provided for the entire batch in GB.
@@
a During inference time measurements with the MACE source code, we were not able to reproduce the original results [33]. Thus, we re-run MACE experiments using the same training setup as for ICTP; see Appendix D.2.
@@
Extrapolation to out-of-domain data. We further assess the performance of ICTP models using the 3BPA data set [48]. The training data set comprises 500 configurations, total energies, and atomic forces acquired from molecular dynamics at 300 K. The test data set is obtained from simulations at 300 K, 600 K, and 1200 K. We also test our models using energies and forces along dihedral rotations of the molecule. Table 2 shows that ICTP models trained using 450 configurations perform on par with state-of-the-art spherical models, similar to the results for rMD17. However, we were not able to reproduce the original results using the current MACE source code and the described training setup [33]. Therefore, for a fair comparison, we unified the training setup for ICTP and MACE (see Appendix D.2) and Table 2 reports the corresponding results for 450 training configurations. In Table A1, we present the results obtained for ICTP and MACE trained using 50 configurations.
@@
Table 3: Energy (E) and force (F) root-mean-square errors (RMSEs) for the acetylacetone data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, while the remaining 50 are used for early stopping. All values for ICTP models are obtained by averaging over five independent runs. For MACE and NequIP, the results are reported for three independent runs. The standard deviation is provided if it is available. Best performances, considering the standard deviation, are highlighted in bold.
@@
a Similar to Table 2, we re-run MACE experiments using the same training setup as for ICTP; see Appendix D.2.
@@ 6 Conclusions and limitations
This work introduces many-body equivariant MPNNs based on higher-rank irreducible Cartesian tensors, offering an alternative to spherical models. We assess the performance of resulting MPNNs using molecular benchmark data sets, such as rMD17, 3BPA, and acetylacetone. In these experiments, MPNNs based on irreducible Cartesian tensors show a lower computational cost of individual operations compared to spherical counterparts. Furthermore, we demonstrate that these Cartesian models achieve accuracy and generalization capability on par with or better than state-of-the-art spherical models while memory consumption is comparable. Our results hold across the typical range of tensor ranks used in modeling many-body interactions and relevant physical properties, i.e., $L\leq 4$.
Limitations. We emphasize our focus on introducing MPNNs based on irreducible Cartesian tensors and prove their equivariance. We adapted the MACE architecture, which uses only tensors of even parity, to enable a fair comparison with spherical models. Further modifications to the architecture are possible, which we expect to exploit the full potential of irreducible Cartesian tensors. Future work should also extend the assessment of MPNNs based on irreducible Cartesian tensors using larger and more diverse atomic systems.
@@
The source code will be made publicly available upon publication.
@@ B.1 Normalization constants
with $L=l_{1}+l_{2}+l_{3}$ and $L_{i}=L-2l_{i}-1$. Here, $C_{l_{1}l_{2}l_{3}}$ is defined such that an $l_{3}$-fold contraction of the tensor $\mathbf{z}_{l_{3}}$, obtained through the irreducible Cartesian tensor product between $\mathbf{x}_{l_{1}}$ and $\mathbf{y}_{l_{2}}$ ($\mathbf{x}_{l_{1}}$ and $\mathbf{y}_{l_{2}}$ are obtained using Eq. (1)), with the unit vector $\hat{\mathbf{r}}$ yields unity. Please refer to the original publication for the motivation behind $D_{l_{1}l_{2}l_{3}}$ [45].
@@ B.2 Further details on the equivariant message passing
where $W_{Z_{u}kL\eta_{\nu}}^{(t)}$ denotes a learnable weight matrix which depends on the chemical element $Z_{u}$ and rank $L$ and which elements are initialized by picking the respective entries from a normal distribution with zero mean and variance of $1/\mathrm{len}(\eta_{\nu})$. The updated node embeddings are further obtained as a linear function of $\big{(}\mathbf{m}_{ukL}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}$ and the residual connection [79, 33]
@@
Full and symmetric product basis. In the MACE architecture, the authors pre-compute products between the generalized Clebsch–Gordan coefficients, which define the interaction between $l_{1}$- and $l_{2}$-rank spherical tensors, and the learnable weight $W_{Z_{u}\eta_{\nu}kL}^{(t)}$ of the linear expansion in Eq. (A1[33]. This approach reduces the computational cost of constructing the product basis with spherical tensors as the effective number of evaluated tensor products is smaller. For irreducible Cartesian tensors, operations like matrix-vector and matrix-matrix products define the interaction between $l_{1}$- and $l_{2}$-rank tensors. Thus, an equivalent operation to those proposed for spherical tensors is generally impossible or would lead to an architecture different from MACE.
Note that one of our main goals is to demonstrate that a many-body equivariant message-passing architecture defined using higher-rank irreducible Cartesian tensors can be as expressive as the one using spherical tensors. Therefore, we compute $\mathbf{m}_{ukL}^{(t)}$ by directly evaluating the $\nu$-fold tensor product, which effectively leads to a larger number of operations than for MACE. We refer to the models based on this architecture as irreducible Cartesian tensor potentials (ICTPs) with the full product basis, or ICTPfull. We also noticed that the number of tensor products, $\mathrm{len}(\eta_{\nu})$, leading to the tensor of rank $L$, can be reduced by the symmetry of the two-fold tensor products, i.e., $\mathbf{A}_{l_{1}}\otimes\mathbf{A}_{l_{2}}=\mathbf{A}_{l_{2}}\otimes\mathbf{A}_{l_{1}}$ if $\mathbf{A}_{l_{1}}$ and $\mathbf{A}_{l_{2}}$ are symmetric. We refer to this design choice as ICTPs with the symmetric product basis or ICTPsym.
Coupled and uncoupled feature channels. Using coupled feature channels instead of uncoupled ones in Eq. (A1) can improve the performance of the final model. Additionally, the number of feature channels and, thus, the overall computational cost can be reduced when constructing the product basis. However, the number of parameters and, thus, the expressiveness of the model can be preserved. Specifically, we can define a linear expansion for combining many-body equivariant features as
@@
where $W_{Z_{u}kk^{\prime}L\eta_{\nu}}^{(t)}$ denotes a learnable weight matrix which depends on the chemical element $Z_{u}$ and rank $L$ and which elements are initialized by picking the respective entries from a normal distribution with zero mean and variance of $1/(\sqrt{d_{t}}\times\mathrm{len}(\eta_{\nu}))$. The construction of the product basis in the latent feature space requires encoding the two-body features $\mathbf{A}_{ukl_{\xi}}^{(t)}$ using the learnable weight matrices in Eq. (6). The linear expansion is then constructed in the latent feature space and decoded using the learnable weight matrices of the linear function in Eq. (A2). We refer to this design choice as ICTPs with the product basis constructed in the latent feature space or ICTPlt. Combining it with the symmetric product basis, we obtain ICTPsym+lt.
@@
B.3 Asymptotic computational complexity

### v2

@@ 4.2 Equivariant message-passing based on irreducible Cartesian tensors
The following section introduces the basic operations for constructing equivariant MPNNs based on irreducible Cartesian tensors. Using their irreducible tensor products, we demonstrate how to build equivariant two- and many-body features, crucial for modeling many-body interactions in molecular and materials systems. Particularly, we focus on MLIPs based on equivariant MPNNs and extend the state-of-the-art MACE architecture [33] to the Cartesian basis. Following the MACE architecture, we use only even tensor products. We split vectors $\mathbf{r}_{uv}=\mathbf{r}_{u}-\mathbf{r}_{v}\in\mathbb{R}^{3}$ from atom $u$ to atom $v$, schematically shown in Fig. 1 (a), into their radial and angular components (unit vectors), i.e., $r_{uv}=\lVert\mathbf{r}_{uv}\rVert_{2}\in\mathbb{R}$ and $\hat{\mathbf{r}}_{uv}=\mathbf{r}_{uv}/r_{uv}\in\mathbb{R}^{3}$, respectively. In the $t$-th message-passing layer, edges $\{u,v\}$ are embedded using a fully connected neural network $R_{kl_{1}l_{2}l_{3}}^{(t)}:\mathbb{R}\rightarrow\mathbb{R}$ with $k$ output feature channels. The radial function $R_{kl_{1}l_{2}l_{3}}^{(t)}$ takes as an input radial distances $r_{uv}$, which are embedded through Bessel functions and multiplied by a smooth polynomial cutoff function [60]. Finally, we use irreducible Cartesian tensors $\mathbf{T}_{l}(\hat{\mathbf{r}})$, similar to spherical tensors $\mathbf{Y}^{l}(\hat{\mathbf{r}})$, to embed unit vectors into the tensor space of maximal rank $l_{\mathrm{max}}$.
@@
Equivariant many-body features. The importance of many-body terms arises from the fact that the interaction between atoms changes when additional atoms are present; see Appendix A. Furthermore, many-body terms are often required to ensure the generalization of interatomic potentials, i.e., their ability to accurately predict energies and forces for temperatures and stoichiometries on which they were not trained [71]. Here, we construct $(\nu+1)$-body equivariant features from $\big{(}\mathbf{A}_{ukl_{\xi}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{\xi}}}$ obtained using Eqs. (4) or (5). The $\nu$-fold Cartesian tensor product, which yields $(\nu+1)$-body features represented by an irreducible Cartesian tensor of rank $L$, reads
@@
Runtime considerations. When choosing an architecture to implement MLIPs, the runtime per energy and force evaluation is crucial. The computational complexity as a function of rank $L$ is $\mathcal{O}\left(9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$ for the irreducible Cartesian tensor product and $\mathcal{O}\left(L^{5}\right)$ for the Clebsch–Gordan one; see Appendix B.3 for more details. Thus, equivariant convolutions based on spherical tensors are more computationally efficient for $L\rightarrow\infty$ than those based on irreducible Cartesian tensors. However, state-of-the-art models and physical properties typically require $L\leq 4$ [33, 72], making sub-leading terms and implementation-dependent amplitudes crucial. Based on our analysis, for $L\leq 4$, we can expect equivariant convolutions based on irreducible Cartesian tensors to be more computationally and memory efficient than their spherical counterparts. For $L\leq 4$, the $\nu$-fold tensor product in the Cartesian basis can also offer computational advantages. Its cost, as a function of rank $L$ and correlation order $\nu$, is $\mathcal{K}(9^{L}L!/(2^{L/2}(L/2)!))^{\nu-1}$ and $\mathcal{K}L^{\frac{1}{2}\nu(\nu+3)}$ for ICTP and MACE, respectively. The pre-factor $\mathcal{K}$, which counts all possible $\nu$-fold tensor products, is removed in MACE through generalized Clebsch-Gordan coefficients, though these coefficients increase the memory spherical models use. Therefore, it is essential to consider the inference times for a fair comparison, which we provide in Section 5.
Equivariance and tracelss property of message-passing layers. We conclude this section by giving a theoretical result that ensures the equivariance of message-passing layers based on irreducible Cartesian tensors and their irreducible tensor products to actions of the orthogonal group. The proof is provided in Appendix C. We also prove in Appendix D that these message-passing layers preserve the traceless property of irreducible Cartesian tensors.
@@
Proposition 4.2.
The message-passing layers based on irreducible Cartesian tensors and their irreducible tensor products preserve the traceless property of irreducible Cartesian tensors.
@@ 5 Experiments and results
This section presents the results for the five benchmark data sets: rMD17, MD22, 3BPA, acetylacetone, and Ta–V–Cr–W. We describe data sets and training details in Appendices E.1 and E.2, respectively.
Figure 2: Inference times and memory consumption as a function of the tensor rank $L$ (a)–(b) and the correlation order $\nu$ (c)–(d). All results are obtained for the 3BPA data set and $l_{\mathrm{max}}=L$. We used eight feature channels to allow experiments with larger $\nu$ values. MACE models use intermediate tensors with $l>l_{\mathrm{max}}$ for their product basis, which we fixed to $l=l_{\mathrm{max}}$. Otherwise, pre-computing generalized Clebsch–Gordan coefficients for $\nu>4$ would require more than 2 TB of RAM. For ICTP, we used the full product basis to compute the same number of $\nu$-fold tensor products as in MACE.
Scaling and computational cost. The expressive power and computational efficiency of equivariant many-body message-passing potentials depend on the tensor ranks employed in equivariant message passing and embedding local atomic environments, i.e., $L$ and $l_{\mathrm{max}}$, respectively, as well as the correlation order $\nu$. Recent work has shown that for identifying environments with $L$-fold symmetries, at least rank-$L$ tensors are required [73]. These symmetries are typically lifted in atomistic simulations, motivating the use of $L\leq 4$. Higher body-order correlations $\nu$, in turn, are required if atomic environments are degenerate to a lower body-order correlation $\nu-1$ [74, 73]. Figure 2 demonstrates inference times and memory consumption of models based on irreducible Cartesian and spherical tensors, i.e., ICTP and MACE, respectively, as a function of the tensor rank and the correlation order. Table A1 presents the corresponding numerical results. We find that irreducible Cartesian tensors outperform spherical ones for most parameter values. In particular, irreducible Cartesian tensors allow spanning the $\nu$-space more efficiently, in line with our theoretical results in Section 4.2.
Table 1: Energy (E) and force (F) mean absolute errors (MAEs) for the rMD17 data set. E- and F-MAE are given in meV and meV/Å, respectively. Results are shown for models trained using $N_{\text{train}}=\{950,50\}$ configurations randomly drawn from the data set, with further 50 used for early stopping. All values are obtained by averaging over five independent runs, with the standard deviation provided if available. Best performances, considering the standard deviation, are highlighted in bold.
@@
We further evaluate the performance of ICTP using the MD22 data set, which contains seven molecular systems with 42–370 atoms [48]. This data set spans four major classes of biomolecules and supramolecules and was designed to challenge short-range models. Table A2 shows that ICTP achieves an accuracy on par with or better than state-of-the-art models, including long-range ones.
Table 2: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set. E- and F-RMSE are given in meV and meV/Å, respectively. Results are shown for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, with further 50 used for early stopping. All ICTP results are obtained by averaging over five independent runs. For MACE and NequIP, the results are reported for three runs. The standard deviation is provided if it is available. Best performances, considering the standard deviation, are highlighted in bold. Inference time and memory consumption are measured for a batch size of 100. Inference time is reported per structure in ms, while memory consumption is provided for the entire batch in GB.
@@
a During inference time measurements with the MACE source code, we were not able to reproduce the original results [33]. Thus, we re-run MACE experiments using a training setup similar to that of ICTP; see Appendix E.2.
@@
Extrapolation to out-of-domain data. We continue to assess the performance of ICTP models using the 3BPA data set [49]. The training data set comprises 500 configurations, total energies, and atomic forces acquired from molecular dynamics at 300 K. The test data set is obtained from simulations at 300 K, 600 K, and 1200 K. We also test our models using energies and forces along dihedral rotations of the molecule. Table 2 shows that ICTP models trained using 450 configurations perform on par with state-of-the-art spherical models, similar to the results for rMD17. However, we were not able to reproduce the original results using the current MACE source code and the described training setup [33]. Therefore, for a fair comparison, we unified the training setup for ICTP and MACE (see Appendix E.2) and Table 2 reports the corresponding results for 450 training configurations. In Table A3, we present the results obtained for ICTP and MACE trained using 50 configurations.
[…]
@@
Table 3: Energy (E) and force (F) root-mean-square errors (RMSEs) for the acetylacetone data set. E- and F-RMSE are given in meV and meV/Å, respectively. Results are shown for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, with further 50 used for early stopping. All ICTP results are obtained by averaging over five independent runs. For MACE and NequIP, the results are reported for three runs. The standard deviation is provided if it is available. Best performances, considering the standard deviation, are highlighted in bold.
@@
a Similar to Table 2, we re-run MACE experiments using the similar training setup as for ICTP; see Appendix E.2.
@@
Multicomponent alloys. We finally evaluate the ICTP and MACE models using the Ta–V–Cr–W data set, designed to assess the performance of state-of-the-art MLIPs in modeling chemically complex multicomponent systems. In this evaluation, we attempt to predict energies and forces for Ta–V–Cr–W subsystems under two scenarios: The 0 K energies and forces in binary, ternary, and quaternary systems and near-melting temperature energies and forces in 4-component disordered alloys. Table A6 shows that ICTPsym outperforms MACE in nearly all subsystems, particularly in energy prediction. ICTP achieves an overall accuracy of 1.38 $\pm$ 0.09 meV/atom for energies and 0.028 $\pm$ 0.001 eV/Å for forces, compared to 2.19 $\pm$ 0.31 meV/atom and 0.029 $\pm$ 0.001 eV/Å by MACE. However, the $\mathcal{K}$ pre-factor from the $\nu$-fold tensor product results in longer inference times for ICTP than MACE, in line with the discussion for 3BPA.
@@ 6 Conclusions and limitations
This work introduces many-body equivariant MPNNs based on higher-rank irreducible Cartesian tensors, offering an alternative to spherical models and addressing the limitations of state-of-the-art Cartesian models. We assess the performance of resulting MPNNs using five benchmark data sets, such as rMD17, MD22, 3BPA, acetylacetone, and Ta–V–Cr–W. In these experiments, MPNNs based on irreducible Cartesian tensors show a lower computational cost of individual operations compared to spherical counterparts. Furthermore, we demonstrate that these Cartesian models achieve accuracy and generalization capability on par with or better than state-of-the-art spherical models while memory consumption is comparable. Our results hold across the typical range of tensor ranks used in modeling many-body interactions and relevant physical properties, i.e., $L\leq 4$.
Limitations. We emphasize our focus on introducing MPNNs based on irreducible Cartesian tensors and prove their equivariance and traceless property. We adapted the MACE architecture, which uses only even tensor products, to enable a fair comparison with state-of-the-art spherical models. Further modifications to the architecture are possible and necessary, e.g., to reduce the pre-factor arising from the Cartesian product basis, before we can fully exploit the potential of irreducible Cartesian tensors.
Data availability
All data sets used in this study are publicly available: rMD17 (https://doi.org/10.6084/m9.figshare.12672038.v3), MD22 (http://www.sgdml.org), 3BPA (https://github.com/davkovacs/BOTNet-datasets), acetylacetone (https://github.com/davkovacs/BOTNet-datasets), and Ta–V–Cr–W (https://doi.org/10.18419/darus-3516).
@@
The source code is available on GitHub and can be accessed via this link: https://github.com/nec-research/ictp.
@@ Appendix A Background
Higher-body-order local descriptors. Recent advances in MLIPs have been influenced by moment tensor potentials (MTPs) [65] and atomic cluster expansion (ACE) [40]. These approaches enable systematic construction of higher-body-order polynomial basis functions, encompassing representations like atom-centered symmetry functions (ACSFs) [68, 75], smooth overlap of atomic positions (SOAP) [76, 77], Gaussian moments [66, 67], and embedded atom/multi-scale embedded atom method (EAM/MEAM) potentials [78, 79]. Furthermore, a reducible Cartesian tensor can be represented as a linear combination of irreducible spherical counterparts, and vice versa [40, 51]. More general expressions, including tensor contractions, have also been provided, hinting at the relationship between Cartesian and spherical models [80, 81, 82]. Despite the success of MTP and ACE, defining smaller cutoff radii and rigid architecture can result in limited accuracy compared to MPNNs.
@@ B.1 Normalization constants
with $L=l_{1}+l_{2}+l_{3}$ and $L_{i}=L-2l_{i}-1$. Here, $C_{l_{1}l_{2}l_{3}}$ is defined such that an $l_{3}$-fold contraction of the tensor $\mathbf{z}_{l_{3}}$, obtained through the irreducible Cartesian tensor product between $\mathbf{x}_{l_{1}}$ and $\mathbf{y}_{l_{2}}$ ($\mathbf{x}_{l_{1}}$ and $\mathbf{y}_{l_{2}}$ are obtained using Eq. (1)), with the unit vector $\hat{\mathbf{r}}$ yields unity. We refer to the original publication for the motivation behind $D_{l_{1}l_{2}l_{3}}$ [45].
@@ B.2 Further details on the equivariant message passing
where $W_{Z_{u}\eta_{\nu}kL}^{(t)}$ denotes a learnable weight matrix which depends on the chemical element $Z_{u}$ and rank $L$ and which elements are initialized by picking the respective entries from a normal distribution with zero mean and a standard deviation of $1/\mathrm{len}(\eta_{\nu})$. The updated node embeddings are further obtained as a linear function of $\big{(}\mathbf{m}_{ukL}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}$ and the residual connection [83, 33]
@@
Full and symmetric product basis. In the MACE architecture, the authors pre-compute products between the generalized Clebsch–Gordan coefficients, which define the interactions of $\{l_{1},\cdots,l_{\nu}\}$-rank spherical tensors, and the learnable weight $W_{Z_{u}\eta_{\nu}kL}^{(t)}$ of the linear expansion in Eq. (A1[33]. This approach reduces the computational cost of constructing the product basis with spherical tensors as the effective number of evaluated tensor products is smaller by $\mathcal{K}=\mathrm{len}\left(\eta_{\nu}\right)$. For irreducible Cartesian tensors, operations like matrix-vector and matrix-matrix products define the interactions between $\{l_{1},\cdots,l_{\nu}\}$-rank tensors. Thus, an equivalent operation to those proposed for spherical tensors may be generally impossible for irreducible Cartesian tensors or would lead to an architecture different from MACE.
Note that one of our main goals is to demonstrate that a many-body equivariant message-passing architecture defined using higher-rank irreducible Cartesian tensors can be as expressive as the one using spherical tensors. Therefore, we compute $\mathbf{m}_{ukL}^{(t)}$ by iteratively evaluating all possible $\nu$-fold tensor products, which effectively leads to a larger number of operations than for MACE. We refer to the models based on this architecture as irreducible Cartesian tensor potentials (ICTPs) with the full product basis, or ICTPfull. We also noticed that the number of tensor products $\mathrm{len}(\eta_{\nu})$ leading to the tensor of rank $L$, can be reduced by the symmetry of the two-fold tensor product in Eq. (6), i.e., $\mathbf{A}_{l_{1}}\otimes\mathbf{A}_{l_{2}}=\mathbf{A}_{l_{2}}\otimes\mathbf{A}_{l_{1}}$ if $\mathbf{A}_{l_{1}}$ and $\mathbf{A}_{l_{2}}$ are symmetric. We refer to this design choice as ICTPs with the symmetric product basis or ICTPsym.
Coupled feature channels. Using coupled feature channels instead of uncoupled ones in Eq. (A1) can improve the performance of the final model. Additionally, the number of feature channels and, thus, the overall computational cost can be reduced when constructing the product basis. However, the number of parameters and, thus, the expressive power of the model can be preserved. Specifically, we can define a linear expansion for combining many-body equivariant features as
@@
where $W_{Z_{u}\eta_{\nu}kk^{\prime}L}^{(t)}$ denotes a learnable weight matrix which depends on the chemical element $Z_{u}$ and rank $L$ and which elements are initialized by picking the respective entries from a normal distribution with zero mean and a standard deviation of $1/(\sqrt{d_{t}}\times\mathrm{len}(\eta_{\nu}))$. The construction of the product basis in the latent feature space requires encoding the two-body features $\mathbf{A}_{ukl_{\xi}}^{(t)}$ using the learnable weight matrices in Eq. (6). The linear expansion is then constructed in the latent feature space and decoded using the learnable weight matrices of the linear function in Eq. (A2). We refer to this design choice as ICTPs with the product basis constructed in the latent feature space or ICTPlt. Combining it with the symmetric product basis, we obtain ICTPsym+lt.
@@
B.3 Computational complexity and runtime analysis